S=ut+1/2at^2. Use this formula to find t when s=6 u=11 a=4
Answer by Cromlix(4381)

Hi there,
S = ut + 1/2at^2
6 = 11t + 1/2*4t^2
6 = 11t + 2t^2
Rearrange into quadratic
2t^2 + 11t - 6 = 0
(2t - 1)(t + 6) = 0
t + 6 = 0
t = -6
or
2t - 1 = 0
t = 1/2
So, t can equal -6 or 1/2
Hope this helps:-)